July 29, 2014 BCAL Commissioning
Video Conferencing Information
Meeting Time: 2:15 p.m. Eastern
- To join via a Web Browser, go to the page  https://bluejeans.com/907185247.
- To join via Polycom room system go to the IP Address: 188.8.131.52 (bjn.vc) and enter the meeting ID: 907185247.
- To join via phone, use one of the following numbers and the Conference ID: 907185247.
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
- More information on connecting to bluejeans is available.
Participant Direct Lines
- JLab Phone: in CC F326 is 757-269-6460 (usual room)
- JLab Phone in CC L207 is 757-269-7084
- Phone in the Regina Video-conference Suite is 306-585-4204
- Athens Phone: in Christina's office is 011-30-210-727-6947
- xFig drawing of trigger scintillator and BCAL geometry for cosmic tests (Shayne, Laura).
- Combined USM+JLab parameter plots.
- ccdb discussion
- Plan to run with CDC and magnet.
- Action Items
- Any other business
Attendees: Elton, Mark D., Will M., Laura, Shayne (JLab), Andrei, Zisis (Regina)
- Elton reported that the magnet is being cooled - it is at 8 K and over next week will go down to 4 K. It will then run at 1200 A initially. All detector groups need to plan running with the magnet. BCAL+CDC using cosmics in a magnetic field is a must. All of BCAL is needed for this, so we will aim for a timesharing mode rather than detector North-South sharing mode with the DAQ effort. Groups report to Yi at this time.
- Action Items
- Shayne: Cosmics are being collected mainly from the North (South is occupied by Sascha). The geometrical asymmetry in the top trigger counter was measured to be 14" biases towards the North side. Still, this probably does not explain the difference in the observed rates. Shayne is working on a toy MC to check the rates in the BCAL modules. He has the cosmics angular profile built in.
- Laura: is working on a Table to separate LED and cosmics runs and their stats for easy viewing. Andrei asked for ROOT files that show module, row, column (translated from crate, slot, etc). Mark said that this can be produced. Laura is also working on a toy MC to extract energy depositions for up:downstream.
- Will: is looking at different ADC modes (raw 2 microsec, pulse time, pulse integral). LEDs are run at 6.5-7.0 V that give large enough amplitudes to see the oscillation (remnant of original problem during 2012 beam test albeit much reduced; present also in cosmics but extremely hard to see due to low signal amplitude). Will is using the first two peaks to test his scripts for the Michel electron analysis and muon lifetime extraction. Mark and Will will discuss using higher level objects, correcting for attenuation, and using the sums or geometrical means from up and down sides.
- Andrei reported briefly on the Geant3 simulations. The muons give a well defined peak, although there is a large exponential tail rising up to zero on the x-axis, so fitting needs care. We are looking at peak values right now, means later on.
- System stability is fine, including the chillers.
- Next steps:
- Activate TDCs. DAQ group (Sergey Furlatov) made changes and think we can read them out. Mark said that we could do this as early as this afternoon. Mark strongly recommends we use Orlando's script for the LED data, which turns on 1-string of LEDs on at a time to reduce cross talk that affects the timing signal (leading to incorrect z-position of hit).
- Take more low temp data with LEDs. Look also at cosmics at +12C. There are no plans to take +5C data before Oct-Nov engineering run. This step (+12C) is in the project and needs to be checked off.
- Any other business
- Laura and Shayne will produce reports on their MC work and their code will be checked into the repository for future use.